Math 544 (Section 001)  University of South Carolina
D. Meade  Fall 1996

Day One Questionnaire – Math 544 (Spring 1996)

• Personal Information Name Phone Number E-mail Address Major Year

• \underline{Schedule}

Work (hrs/wk) Relax (hrs/wk) Study (hrs/wk) Course (hrs/wk) Other Courses

• Mathematical Background

Course  When  Where  Who  Grade

Math 142 (or equiv)

Math 241 (or equiv)

Other Post-Calculus Math • Other Information

1. Why are you taking this course?

2. Do you have any specific goals or objectives that you hope to fulfill by taking this course?

3. What is your biggest concern about this course?

4. Do you plan to utilize office hours? If “yes”, what are the best times I could have office hours? I’ll try to accommodate as many people as possible, but it is almost impossible to satisfy all requests.^{1}

1This statement can be phrased in the language of linear algebra. See if you recognize this when we encounter it in Chapter 1.